calculate the minimum thickness needed for an antireflective coating (n=1.39) applied to a glass lens in order to eliminate blue (470nm) reflections for light at normal incidence.

calculate the minimum thickness needed for an antireflective coating (n=1.39) applied to a glass lens inorder to eliminate red (695 nm) reflections for light at normal incidence.
Answer #1



1) lmin = 470 *10^-9/4*1.39 = 8.45 *10^-8 m



2) lmin = 695 *10^-9/4*1.39 = 1.25 *10^-7 m
Answer #2

The minimum thickness needed for an anti reflective coating is

\(t_{\min }=\frac{\lambda}{4 n}\)

forblue

\(\begin{aligned} t_{\min } &=\frac{470 \mathrm{~nm}}{4(1.39)} \\ &=84.53 \mathrm{~nm} \end{aligned}\)

for red

\(t_{\min }=\frac{695 \mathrm{~nm}}{4(1.39)}\)

\(=125 \mathrm{~nm}\)